### v1

@@ Abstract
Recent advancements in Artificial Intelligence (AI) demonstrate significant potential to revolutionize weather forecasting. However, most AI-driven models rely on Numerical Weather Prediction (NWP) systems for initial condition preparation, which often consumes hours on supercomputers. Here we introduce XiChen, the first observation-scalable fully AI-driven global weather forecasting system, whose entire pipeline, from Data Assimilation (DA) to medium-range forecasting, can be accomplished within only 17 seconds. XiChen is built upon a foundation model that is pre-trained for weather forecasting. Meanwhile, this model is subsequently fine-tuned to serve as both observation operators and DA models, thereby scalably assimilating conventional and raw satellite observations. Furthermore, the integration of four-dimensional variational knowledge ensures that XiChen’s DA and medium-range forecasting accuracy rivals that of operational NWP systems, amazingly achieving a skillful forecasting lead time exceeding 8.25 days. These findings demonstrate that XiChen holds strong potential toward fully AI-driven weather forecasting independent of NWP systems.
@@ 1 Introduction
A century ago, Abbe [1] and Bjerknes [2] conceptualized weather forecasting as an initial value problem, where forecasts could be derived by starting from initial conditions estimated from observations and applying physical laws. At that time, atmospheric observations were extremely limited, computers did not exist, and the predictability of weather remained largely unknown. Thus, their proposition was remarkably ambitious [3]. But today, advancements in our understanding of atmospheric mechanisms, the continuous accumulation of observational data, and significant improvements in computational power have collectively enhanced the accuracy of Numerical Weather Prediction (NWP), making it an indispensable part of daily life [4].
NWP consists of two core components. The first is Data Assimilation (DA), which estimates the most likely state of the atmosphere based on recent observational data from satellites, meteorological stations, ships, and other sources to provide accurate initial conditions for forecasting [3, 4]. The second component is the forecast model, which traditionally simulates the spatial-temporal evolution of atmospheric variables by solving a set of complex Partial Differential Equations (PDEs) [3, 4]. Although NWP has undergone a technological revolution [3], several challenges have emerged. The deceleration of Moore’s Law, increasing complexity of modern NWP systems, and substantial improvements in resolution have made it difficult to deliver timely forecasting results within limited computational time [4, 5]. These limitations consequently impede further progress in NWP.
@@
In recent years, the rapid development of Artificial Intelligence (AI) technology has triggered a transformative revolution in weather forecasting [6]. On dedicated computing hardware, such as Graph Processing Units (GPUs), AI-based models—such as FourCastNet [7], Pangu-Weather [5], GraphCast [8], FengWu [9], Fuxi [10], and AIFS [11]—achieved acceleration rates three to four orders of magnitude faster than operational NWP systems while maintaining comparable forecasting performance. However, these AI models still rely on the (re)analysis fields produced by NWP systems to generate forecasts and are far from a fully AI-driven global weather forecasting system [12, 13, 14]. For instance, the analysis fields used by current AI-based forecasting models (e.g., ECMWF’s AIFS [11]) are products generated through the Four-Dimensional Variational (4DVar) method [15]. The 4DVar method employs optimal control theory to assimilate over 10 million observational data points within each 12-hour Data Assimilation Window (DAW) [16]. Notably, it effectively integrates indirect satellite observations, addressing gaps in atmospheric field information that are challenging to capture using sparse conventional observations [16]. Nevertheless, the 4DVar method in operational NWP systems requires hours of computation on supercomputers [17], restricting the volume of observations that can be assimilated within a limited time frame, with only 5-10% of the available satellite data being used [3]. Additionally, 4DVar requires meticulous tuning and precise formulation of a complex background error covariance matrix [18]. It also depends on sophisticated observation operators [19], tangent linear models, and adjoint models [18], all of which demand extensive development time and expertise from highly trained specialists.
@@
In this era of rapid advancements in AI, a bold proposition has emerged: the development of data-driven systems based on AI technology to forecast future weather states [20, 21] without dependence upon the initial fields from NWP systems. Notably, models such as Aardvark [22] and GraphDOP [23] have successfully generated forecasts directly from raw observations, demonstrating promising performance. However, the natural sparsity of observations poses significant challenges to achieving forecast accuracy comparable to operational NWP systems. Moreover, directly taking all observations as inputs without rigorous quality control may compromise forecast accuracy. Recent studies have also explored integrating pre-trained AI-based weather forecasting models into traditional 4DVar frameworks by leveraging automatic differentiation to create hybrid 4DVar systems [24, 25, 26]. Nevertheless, these efforts often rely on idealized models or simulated observations. Even when real-world conventional observations are employed, these approaches still overlook satellite observations, which are critical for operational NWP systems. On the other hand, some researchers have also ambitiously pursued the development of data-driven DA systems [27, 28, 29, 30, 31], such as FengWu-Adas [32], FuXi-DA [13], DiffDA [30], and 4DVarFormer [14, 33]. These models utilize modern neural network architectures to assimilate conventional and satellite observations, integrating them with AI-based weather forecasting models to create fully AI-driven global weather forecasting systems. Despite these advancements, they can only assimilate observation types encountered during training, necessitating retraining and architectural refinement of the entire model to incorporate new observation types. This limitation restricts their scalability compared to traditional DA systems. Consequently, developing fully AI-driven global weather forecasting systems suitable for operational implementation remains an unresolved challenge.
@@
Here, we propose XiChen (Chinese: {CJK}UTF8gbsn羲辰), the first observation-scalable fully AI-driven global weather forecasting system. XiChen is constructed by a foundation model [34] trained on diverse weather data to replace core components of the NWP pipeline. The foundation model is initially pre-trained on ERA5 [35] data to learn fundamental atmospheric dynamical representations as the medium-range weather forecasting model. Subsequently, it is fine-tuned to incorporate the observation operators and the DA capabilities. By incorporating the 4DVar knowledge into the DA model, specifically by leveraging the gradient of the AI-driven 4DVar cost function as model inputs, we propose a cascaded sequential DA framework to effectively assimilate both conventional and raw satellite observations for initializing accurate global weather forecasting. This cascaded DA framework shows the potential to address the compatibility and robustness challenges associated with the future integration of new observations by fine-tuning. In our experiments, XiChen assimilates conventional observations from the Global Data Assimilation System (GDAS) and raw satellite observational data from the Advanced Microwave Sounding Unit-A (AMSU-A) and the Microwave Humidity Sensor (MHS) within 5 seconds. Although XiChen operates on a global grid with a resolution of 1.40625, its accuracy is comparable to that of operational NWP systems across multiple variables in DA and medium-range weather forecasting tasks. Furthermore, the entire pipeline of XiChen, from DA to 10-day medium-range forecasting, can be accomplished within only 17 seconds using a single A100 GPU, achieving over 400-fold acceleration vs the operational NWP system [17]. Notably, XiChen achieves a skillful forecast lead time of over 8.25 days, which is comparable to the Global Forecasting System (GFS) while surpassing the State-Of-The-Art (SOTA) end-to-end global weather forecasting systems: Aardvark [22] (less than 8 days) and GraphDOP [23] (about 5 days). Furthermore, we propose a dual DA framework to circumvent the multi-hour latency characteristic of the Integrated Forecasting System (IFS HRES), thereby enabling real-time acquisition of medium-range forecast products. We comprehensively evaluate various DA observation configurations and their impact on medium-range forecasting, the effect of 4DVar knowledge on DA and forecasting, and XiChen’s performance in the Tropical Cyclone (TC) track task. These findings highlight that XiChen offers a novel avenue for advancing fully AI-driven global weather forecasting through fine-tuning to scalably integrate new observations in the future.
@@ 2 Results
Our experiments evaluated two key aspects of XiChen’s performance: (1) the accuracy and stability of its one-year DA cycle, and (2) the skill of its 10-day medium-range weather forecasts. The DA process employed a cascade framework, sequentially assimilating GDAS prepbufr data, followed by AMSU-A and MHS observations. For the DA cycle experiment, XiChen was initialized at 00:00 UTC on 1 January 2023. The background field was generated by initializing the XiChen forecast model with ERA5 reanalysis data from 00:00 UTC on 27 December 2022, followed by a 5-step rollout with a 24-hour lead time. This forecast field served as the initial condition for the DA cycle, which ran for one year with assimilations at 00:00 and 12:00 UTC daily. We employ a 12-hour DAW wherein observations at 0, 3, 6, and 9 hours after the analysis time are assimilated to initialize the forecasting model for generating the background field for the subsequent DA cycle. This computational process is executed within 5 seconds utilizing a single A100 GPU. Furthermore, we implement a 3-hour DAW that exclusively assimilates observations at the analysis time point to establish the initial conditions for medium-range forecasting. This dual DA framework ensures the fidelity of the assimilation process and for medium-range predictions. Furthermore, XiChen circumvents the multi-hour latency characteristic of IFS, thereby enabling real-time acquisition of medium-range forecast products. Further details of the cascade process and DA cycle are illustrated in Figures S1(a) and S1(b). The weather forecasting model, DA models, and observation operators share the same underlying model architecture and are trained using different data. Details regarding this architecture are provided in Section 4.
For medium-range weather forecasting, we adopted configurations and evaluation metrics from WeatherBench [36] and DABench [33] to generate initial fields and evaluate XiChen’s 10-day forecasts. For details, please see Section 4.8 and Figure S1(c). XiChen’s performance was compared against SOTA operational NWP systems: IFS HRES and GFS. We assessed eight ECMWF Scorecard variables: geopotential at 500 hPa (Z500) and 850 hPa (Z850), 2-meter temperature (T2M), temperature at 850 hPa (T850), Mean Sea Level Pressure (MSLP), 10-meter zonal (10U) and meridional (10V) winds, and specific humidity at 500 hPa (Q500).
Additionally, we evaluated XiChen’s TC assimilation and forecasting capabilities. TC tracks were identified using the TempestExtremes [37] tracker, configured as in NeuralGCM [38], and compared against the International Best Track Archive for Climate Stewardship (IBTrACS) and ERA5. Tracker parameters are detailed in Table S7, and TC track results from forecasts initialized with different observations are shown in Figures S10 to S13. Finally, we present visualizations of the forecasting results produced by XiChen, IFS HRES, and GFS in Figures S14 to S19.
@@ 2.1 One year DA Cycle
The time series of the globally averaged, annual latitude-weighted Root Mean Square Error (RMSE) and Bias for the DA cycle results of XiChen, compared to the operational analysis fields from the GFS and the IFS HRES, are presented in Figure 1. Each subplot corresponds to a variable, and skill (y-axis) is plotted at 12-hour intervals across the 365-day horizons (x-axis). Rows 1 and 3 show RMSE, while rows 2 and 4 show Bias. The original data is represented with reduced opacity to enhance clarity, while solid lines depict values smoothed using an Exponential Moving Average (EMA) with a 15-point window. XiChen demonstrates stable operation over long-term DA cycles, with the RMSE of key variables such as Z500, T850, Z850, T2M, U10, V10, and MSLP outperforming those of GFS and approaching those of IFS HRES. Furthermore, the Bias of XiChen remains closer to zero than GFS throughout the year, highlighting its capability to mitigate the bias drift of the forecasting fields. Since the initial background field is a 5-day forecast initialized using ERA5, there is a significant difference compared to the reference. XiChen achieves stable analysis performance after a spin-up period, similar to the 4DVar approach, which can be attributed to integrating 4DVar knowledge within the XiChen DA framework.
The occasional increase in RMSE, as shown in Figure 1, is primarily caused by missing observations. However, it is evident that XiChen effectively mitigates this issue through subsequent assimilation processes. The assimilation is performed using a cascading approach. When AMSU-A or MHS observations are unavailable, the XiChen system excludes the assimilation of AMSU-A or MHS data. Despite this, the accuracy of the analysis fields remains stable. These findings indicate that XiChen can potentially develop into an operational system.
Figure 1: Comparison of the average latitude RMSE (first and third rows) and Bias (second and fourth rows) of the one-year DA cycle using XiChen (red lines) as well as the operational analysis fields of GFS (yellow lines) and IFS HRES (blue lines) for 4 surface variables, such as MSLP, T2M, U10, and V10, and 4 upper-air variables, including Z500, T850, Z850, and Q500, using testing data during 2023. All analysis fields are evaluated against the ERA5 reanalysis dataset.
@@ 2.2 Medium-range weather forecasting
The primary metric for evaluating the fully AI-driven global weather forecasting system is its performance in medium-range weather forecasts. In this study, we assess the 6-hour interval medium-range forecasts produced by XiChen. The initial fields for the medium-range forecasting experiment are derived from the analysis fields generated within the DA cycle experiment. The timing of these initial fields is selected based on the setups of WeatherBench [36] and DABench [33]. Figure 2 illustrates the performance of the medium-range forecast initialized by the analysis fields obtained from assimilating GDAS prepbufr observations. Each subplot corresponds to a variable and skill (y-axis) is plotted at 6-hour steps over 10-day horizons (x-axis). Rows 1 and 3 show RMSE, rows 2 and 4 show the Anomaly Correlation Coefficient (ACC). Notably, for key variables such as T2M, T850, MSLP, U10, and V10, XiChen delivers forecasts with lower RMSEs compared to the GFS for nearly the entire 10-day lead time. The RMSEs for all variables approach or even surpass the performance of the IFS HRES as the lead time increases. For Q500, XiChen exceeds the GFS after 2 days and the IFS HRES after 4 days. Although the forecast error for Z500 remains higher than that of both the GFS and the IFS HRES during the first 6 days, XiChen surpasses the GFS after 7 days and outperforms IFS HRES after 8 days.
@@
Figure 2: Comparison of the average latitude RMSE (first and third rows) and ACC (second and fourth rows) of the 10-day medium-range weather forecasting using XiChen (red lines) as well as the operational forecasting results of GFS (yellow lines) and IFS HRES (blue lines) for 4 surface variables, such as MSLP, T2M, U10, and V10, and 4 upper-air variables, including Z500, T850, Z850, and Q500, using testing data during 2023. All analysis fields are evaluated against the ERA5 reanalysis dataset.

### v2

@@ Abstract
Artificial intelligence (AI)–driven models have the potential to revolutionize weather forecasting, but still rely on initial conditions generated by costly Numerical Weather Prediction (NWP) systems. Although recent end-to-end forecasting models attempt to bypass NWP systems, these methods lack scalable assimilation of new types of observational data. Here, we introduce XiChen, an observation-scalable fully AI-driven global weather forecasting system, wherein the entire pipeline, from Data Assimilation (DA) to medium-range forecasting, can be accomplished within only 15 seconds. XiChen is built upon a foundation model that is pre-trained for weather forecasting and subsequently fine-tuned to serve as both observation operators and DA models, thereby enabling the scalable assimilation of conventional and raw satellite observations. Furthermore, the integration of Four-Dimensional Variational (4DVar) knowledge ensures XiChen to achieve DA and medium-range forecasting accuracy comparable to operational NWP systems, with skillful forecasting lead time beyond 8.75 days. A key feature of XiChen is its ability to maintain physical balance constraints during DA, enabling observed variables to correct unobserved ones effectively. In single-point perturbation DA experiments, XiChen exhibits flow-dependent characteristics similar to those of traditional 4DVar systems. These results demonstrate that XiChen holds strong potential for fully AI-driven weather forecasting independent of NWP systems.
@@ 1 Introduction
A century ago, Abbe [abbe1901physical] and Bjerknes [bjerknes1904problem] conceptualized weather forecasting as an initial value problem, where forecasts could be derived by starting from initial conditions estimated from observations and applying physical laws. At that time, atmospheric observations were extremely limited, computers did not exist, and the predictability of weather remained largely unknown. Thus, their proposition was remarkably ambitious [bauer2015quiet]. But today, advancements in our understanding of atmospheric mechanisms, the continuous accumulation of observational data, and significant improvements in computational power have collectively enhanced the accuracy of Numerical Weather Prediction (NWP). NWP consists of two core components. The first is Data Assimilation (DA), the process of integrating observations from satellites, meteorological stations, ships, and other sources to estimate the most likely state of the atmosphere and provide accurate initial conditions for forecasting [bauer2015quiet, gettelman2022future]. The second component is the forecast model, which traditionally simulates the spatial-temporal evolution of atmospheric variables by solving a set of complex Partial Differential Equations (PDEs) [bauer2015quiet, gettelman2022future]. Over the past several decades, NWP has undergone a profound technological transformation [bauer2015quiet], making it an indispensable part of modern society [gettelman2022future].
@@
However, the deceleration of Moore’s Law, the increasing complexity of modern NWP systems, and substantial improvements in resolution have made it difficult to produce timely forecasts within limited computational resources [gettelman2022future, bi2023accurate]. These constraints have consequently hindered further progress in NWP. In recent years, the rapid development of Artificial Intelligence (AI) technology has triggered a transformative revolution in weather forecasting [ling2024improving]. On dedicated computing hardware, such as Graph Processing Units (GPUs), AI-based models—such as FourCastNet [kurth2023fourcastnet], Pangu-Weather [bi2023accurate], GraphCast [lam2023learning], FengWu [ailabchen2025fengwu], Fuxi [fudanchen2023fuxi], and AIFS [lang2024aifs]—achieved acceleration rates three to four orders of magnitude faster than operational NWP systems while maintaining comparable forecasting performance. However, these AI models still rely on the (re)analysis fields produced by NWP systems to generate forecasts and are far from a fully AI-driven global weather forecasting system [chen2023towards, xu2024fuxida, wang2024accurate]. For instance, the analysis fields used by current AI-based forecasting models (e.g., ECMWF’s AIFS [lang2024aifs]) are products generated through the Four-Dimensional Variational (4DVar) method [courtier1994strategy]. The 4DVar method employs optimal control theory to assimilate over 10 million observational data points within each 12-hour Data Assimilation Window (DAW) [lupu2017data]. It effectively integrates indirect satellite observations, addressing gaps in direct atmospheric observations that are challenging to capture using sparse conventional observations [lupu2017data]. Nevertheless, the 4DVar method in operational NWP systems requires hours of computation on supercomputers [bauer2020ecmwf], restricting the volume of observations that can be assimilated within a limited time frame, with only 5-10% of the available satellite data being used [bauer2015quiet]. Additionally, 4DVar requires meticulous tuning and precise formulation of a complex background error covariance matrix [gustafsson2018survey]. It also depends on sophisticated observation operators [liang2023machine], tangent linear models, and adjoint models [gustafsson2018survey], all of which demand extensive development time and expertise from highly trained specialists.
@@
In this era of rapid advancements in AI, a bold proposition has emerged: the development of data-driven systems based on AI technology to forecast future weather states [schultz2021can, mcnally2024data] without dependence upon the initial fields from NWP systems. Notably, models such as Aardvark [allen2025end] and GraphDOP [alexe2024graphdop] have successfully generated forecasts directly from raw observations, demonstrating promising performance. However, the natural sparsity of observations and the lack of rigorous quality control pose significant challenges to achieving forecast accuracy comparable to operational NWP systems. Recent studies have also explored integrating pre-trained AI-based weather forecasting models into traditional 4DVar frameworks by leveraging automatic differentiation to create hybrid 4DVar systems [xiao2024fengwu4dvar, li2024fuxien4dvar, fan2025latent4dvar]. Nevertheless, these efforts often rely on idealized models or simulated observations. Even when real-world conventional observations are employed, these approaches still overlook satellite observations, which are critical for operational NWP systems. On the other hand, some researchers have also ambitiously pursued the development of data-driven DA systems [cintra2016tracking, wu2021fast, arcucci2021deep, huang2024diffda, wang2024four], such as FengWu-Adas [chen2023fengwuadas], FuXi-DA [xu2024fuxida], DiffDA [huang2024diffda], and 4DVarFormer [wang2024accurate, wang2024benchmark]. These models utilize modern neural network architectures to assimilate conventional and satellite observations, integrating them with AI-based weather forecasting models to create fully AI-driven global weather forecasting systems. Despite these advancements, they can only assimilate observation types encountered during training, necessitating retraining and architectural refinement of the entire model to incorporate new observation types. This limitation restricts their scalability compared to traditional DA systems. Consequently, developing fully AI-driven global weather forecasting systems suitable for operational implementation remains an unresolved challenge.
@@
Here, we propose XiChen (Chinese: 羲辰), an observation-scalable fully AI-driven global weather forecasting system. XiChen is constructed by a foundation model [awais2025foundation] trained on diverse weather data to replace core components of the NWP pipeline. The foundation model is initially pre-trained on ERA5 [hersbach2020era5] data to learn fundamental atmospheric dynamical representations, serving as the medium-range weather forecasting model. Subsequently, it is fine-tuned to incorporate the observation operators and the DA capabilities. By incorporating the 4DVar knowledge into the DA model, specifically by leveraging the gradient of the AI-driven 4DVar cost function as model inputs, we propose a cascaded sequential DA framework to effectively assimilate both conventional and raw satellite observations for initializing accurate global weather forecasting. This cascaded DA framework shows the potential to address the compatibility and robustness issues associated with the future integration of new observations by fine-tuning. In our experiments, XiChen assimilates conventional observations from the Global Data Assimilation System (GDAS), raw satellite data from the Microwave Humidity Sensor (MHS) and Advanced Microwave Sounding Unit-A (AMSU-A), as well as wind field retrievals from the Advanced Scatterometer (ASCAT) and satellite-derived winds (SATWND). Although XiChen operates on a global grid with a resolution of 1.40625, its accuracy is comparable to that of operational NWP systems across multiple variables in DA and medium-range weather forecasting tasks. Furthermore, the entire pipeline of XiChen, from DA to 10-day medium-range forecasting, can be accomplished within only 15 seconds using a single A100 GPU, achieving over 400-fold acceleration compared to the operational NWP system [bauer2020ecmwf], with a skillful forecast lead time of over 8.75 days, surpassing the Global Forecasting System (GFS). A distinctive capability of XiChen lies in its ability to maintain physical balance constraints during DA tasks. This enables the system to utilize observed variables to effectively correct other unobserved but interrelated ones, ensuring internal consistency. In single-point perturbation DA experiments, XiChen exhibited flow-dependent characteristics comparable to those observed in the traditional 4DVar system [zhu2022four]. A comprehensive evaluation is presented for three key aspects: 1) the impacts of various DA observation configurations on medium-range forecasting; 2) the effects of 4DVar knowledge on both DA and forecasting; and 3) XiChen’s performance on the Tropical Cyclone (TC) track task. The promising findings highlight that XiChen offers a novel avenue for advancing fully AI-driven global weather forecasting through fine-tuning to scalably integrate new observations in the future.
@@ 2 Results
Our experiments evaluated two key aspects of XiChen’s performance: (1) the accuracy and stability of its one-year DA cycle, and (2) the skill of its 10-day medium-range weather forecasts. The DA process employed a cascade framework, sequentially assimilating GDAS prepbufr data, followed by MHS, ASCAT, AMSU-A, and SATWND observations. For the DA cycle experiment, XiChen was initialized at 00:00 UTC on 1 January 2023. The background field was generated by initializing the XiChen forecast model with ERA5 reanalysis data from 00:00 UTC on 29 December 2022, followed by a 3-step rollout with a 24-hour lead time. This forecast field served as the initial condition for the DA cycle, which ran for one year with assimilations at 00:00 and 12:00 UTC daily. We employ a 12-hour DAW wherein observations at 0, 3, 6, and 9 hours after the analysis time are assimilated to initialize the forecasting model for generating the background field for the subsequent DA cycle. To improve the quality of the analysis field, we constructed an ensemble DA using 11 ensemble members within the 12-hour DAW. The weight of each ensemble member was determined based on its distance from the GDAS prepbufr. For detailed information on the ensemble averaging configuration, refer to Section 4. Furthermore, we implement a 6-hour DAW to establish the initial conditions for medium-range forecasting. This computational process is executed within 3 seconds utilizing a single A100 GPU. This dual DA framework ensures the fidelity of the assimilation process for medium-range predictions. Further details of the cascade process and DA cycle are illustrated in Figures S1(a) and S1(b). The weather forecasting model, DA models, and observation operators share the same underlying model architecture and are trained using different data. Details regarding this architecture are provided in Section 4. For medium-range weather forecasting, we adopted configurations and evaluation metrics from WeatherBench [rasp2020weatherbench] and DABench [wang2024benchmark] to generate initial fields and evaluate XiChen’s 10-day forecasts. For details, please see Section 4.8 and Figure S1(c).
Additionally, we evaluated XiChen’s TC assimilation and forecasting capabilities. TC tracks were identified using the TempestExtremes [ullrich2021tempestextremes] tracker, configured as in NeuralGCM [kochkov2024neural], and compared against the International Best Track Archive for Climate Stewardship (IBTrACS) and ERA5. Tracker parameters are detailed in Table S9, and TC track results from forecasts at different lead times are shown in Figures S9. Finally, we present visualizations of the forecasting results produced by XiChen, IFS HRES, and GFS in Figures S10 to S15.
@@ 2.1 One year DA Cycle
The DA performance of XiChen was systematically evaluated against State-Of-The-Art (SOTA) operational NWP systems, specifically the IFS HRES and GFS. These variables were carefully selected to provide a comprehensive assessment of the system’s predictive capabilities across multiple atmospheric levels and meteorological parameters.
Figure 1 presents the time series of globally averaged annual Root Mean Square Error (RMSE) for the DA cycle results of XiChen, using the ERA5 data as the reference. The RMSE values are compared against those of the operational analysis fields from the GFS and IFS HRES models. Each subplot corresponds to a specific variable, with skill (y-axis) plotted at 12-hour intervals over a 365-day forecast horizon (x-axis). The original data is displayed with reduced opacity for clarity, while solid lines represent values smoothed using an Exponential Moving Average (EMA) with a 15-point window. XiChen demonstrates stable performance throughout the year‑long DA cycle and effectively leverages limited observations to estimate key atmospheric variables. For surface variables with relatively dense direct measurements, such as T2M, U10, and V10, XiChen achieves lower RMSE than GFS. XiChen also attains RMSE values comparable to or lower than those of GFS for Z300, Z500, Z700, and T850. In contrast, XiChen exhibits larger errors for T300, T500, U300, U500, U700, Q300, Q500, and Q700, largely due to the scarcity of observations at upper levels. Occasional increases in RMSE, as seen in Figure 1, are primarily attributed to missing observations. However, XiChen effectively addresses these issues through subsequent assimilation processes.
Figure 1: Comparison of the average latitude RMSE of the one-year DA cycle using XiChen (red lines) as well as the operational analysis fields of GFS (yellow lines) and IFS HRES (blue lines) using testing data during 2023. The analysis encompassed 16 distinct variables, including geopotential heights at 300 hPa (Z300), 500 hPa (Z500), and 700 hPa (Z700); 2-meter temperature (T2M); temperatures at 300 hPa (T300), 500 hPa (T500), 700 hPa (T700), and 850 hPa (T850); u- and v-components of the wind at 10 meters (U10 and V10, respectively); u-components of the wind at 300 hPa (U300), 500 hPa (U500), and 700 hPa (U700); and specific humidity at 300 hPa (Q300), 500 hPa (Q500), and 700 hPa (Q700). All analysis fields are evaluated against the ERA5 reanalysis dataset.
@@ 2.2 Medium-range weather forecasting
The primary criterion for evaluating the fully AI-driven global weather forecasting system lies in its efficacy in generating medium-range global weather forecasts. In this investigation, we evaluate the medium-range global weather forecasts produced at 6-hour intervals by XiChen. The forecast output is systematically compared against reanalysis data from the ERA5 dataset. The temporal initialization of these analysis fields aligns with the configurations established by WeatherBench [rasp2020weatherbench] and DABench [wang2024benchmark], ensuring consistency with prior benchmarks. XiChen’s forecasting performance is evaluated against SOTA operational NWP systems, specifically the IFS HRES and GFS models.
Figure 2 illustrates the averaged latitude-weighted RMSE and Anomaly Correlation Coefficient (ACC), both as a function of forecast lead times over 10 days. Each subplot corresponds to a variable and skill (y-axis) is plotted at 6-hour steps over 10-day horizons (x-axis). Rows 1 and 3 show RMSEs, rows 2 and 4 show the ACCs. When using data from the ERA5 dataset as the initial fields, XiChen outperforms IFS HRES and GFS in all evaluated variables (the red dashed lines). XiChen achieves a skillful weather forecasting lead time exceeding 10.25 days (with ACC of Z500 >0.6>0.6). This performance rivals SOTA medium-range global weather forecasting models trained on 0.25-degree high-resolution datasets—Pangu [bi2023accurate], GraphCast [lam2023learning], FengWu [ailabchen2025fengwu], and FuXi [fudanchen2023fuxi].
@@
Figure 2: Comparison of the average latitude RMSE (first and third rows) and ACC (second and fourth rows) of the 10-day medium-range weather forecasting using XiChen (red lines) as well as the operational forecasting results of GFS (yellow lines) and IFS HRES (blue lines) using testing data during 2023. The evaluation encompasses eight meteorological variables: geopotential height at 500 hPa (Z500) and 850 hPa (Z850), 2-meter temperature (T2M), temperature at 850 hPa (T850), Mean Sea Level Pressure (MSLP), 10-meter zonal (10U) and meridional (10V) wind components, as well as specific humidity at 500 hPa (Q500). All analysis fields are evaluated against the ERA5 reanalysis dataset.